OpenStudy (anonymous):

solve the equations: -5y-9=-(y-1) 3y+20=3+2y

geerky42 (geerky42):

Are you asking two questions? For first equation, distribute -1 on right side. You should have \(-5y-9 = -y+1\) Now add both sides by y: \(-5y-9 \color{red}{+y}= -y+1\color{red}{+y}\) \(-4y-9 = \cancel{-y}+1+\cancel{y}\) \(-4y-9 = 1\) See how I approach this problem? Can you do the rest?

OpenStudy (anonymous):

Yes two different questions